OpenStudy (mchilds15):

How do you find the average rate of change for f(x) from x=4 to x=7?

OpenStudy (jhannybean):

\[\sf \text{avg.rate of change} = \frac{f(b)-f(a)}{b-a}\]

OpenStudy (jhannybean):

You're missing your function though.

OpenStudy (mchilds15):

Do i use vertex or standard?

OpenStudy (jhannybean):

Just standard should be fine.

OpenStudy (jhannybean):

Would you mind posting your entire question up? :)

OpenStudy (mchilds15):

Oh, well then \[2x^2-4+21\] is my equation in standard form

OpenStudy (mchilds15):

I just don't know where to go from there. It's been awhile lol

OpenStudy (jhannybean):

Awesome :) \[f(x)= 2x^2-4x+21\] right?

OpenStudy (camerondoherty):

\[ax^2+bx+c = 0? \]

OpenStudy (camerondoherty):

Identify Parts, sub into equation

OpenStudy (jhannybean):

Think you might be missing an x there.

OpenStudy (mchilds15):

Yes Jhan

OpenStudy (camerondoherty):

Sorry if I barged in btw, counldn't resist *u*

OpenStudy (mchilds15):

Tis ok Cam

OpenStudy (jhannybean):

Alright.\[\frac{f(7)-f(4)}{7-4}\]What is \(f(7) =~?\) What is \(f(4) =~?\)

OpenStudy (mchilds15):

/.\ Ugh I don't know

OpenStudy (jhannybean):

Your equation: \(f(x) = 2x^2-4x+21\) Just plug in 7 and 4 into your \(x\) values.

OpenStudy (mchilds15):

OH... OH OH OH... So i plug them in one at a time and the answer would be the f(x)?

OpenStudy (jhannybean):

The answer will be a quantitative value :)

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @Jhannybean Alright.\[\frac{f(7)-f(4)}{7-4}\]What is \(f(7) =~?\) What is \(f(4) =~?\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (jhannybean):

Answer those two questions and we'll go from there! :)

OpenStudy (camerondoherty):

Hint: To find f(7) and f(4) you plug them into the main equation \[f(x)=2x^2−4x+21 \]\[f(7)=2(7)^2−4(7) +21 \]

OpenStudy (mchilds15):

.... would i multiply the 2*7 first or square 7

OpenStudy (jhannybean):

square 7

OpenStudy (jhannybean):

PEMDAS, remember?

OpenStudy (mchilds15):

Yes lol So f(7)=91 and f(4)=37 ?

OpenStudy (jhannybean):

yep

OpenStudy (camerondoherty):

Yep, Great Job! Now you plug those values into the fraction @Jhannybean provided c: \[\frac{ f(7)-f(4) }{ 7-4 }\]

OpenStudy (jhannybean):

\[\frac{91-37}{7-4}\]Just evaluate that :)

OpenStudy (mchilds15):

Which would be 54/3

OpenStudy (jhannybean):

and what does that reduce to?

OpenStudy (mchilds15):

18?

OpenStudy (jhannybean):

Good job :)

OpenStudy (camerondoherty):

Thats ur rate of change c:

OpenStudy (mchilds15):

xO Mind blown Thanks for teaching me ladies :3

OpenStudy (jhannybean):

no problem :)

OpenStudy (camerondoherty):

Sorry for cutting in cx